OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)?

OpenStudy (anonymous):

First Find Slope

OpenStudy (anonymous):

\[\tfrac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

what's that^

OpenStudy (anonymous):

barely see that^

OpenStudy (anonymous):

then... use....\[y-y1 = m(x-x1)\]

OpenStudy (anonymous):

hold on lemme plug the numbers in.......

OpenStudy (anonymous):

find slope m = \[\frac{ y1-y2 }{ x1-x2 }\]

OpenStudy (anonymous):

-2 - 2^1 = m(0-8^1)

OpenStudy (anonymous):

use any of the point now and find b in the below equation y=mx+b

OpenStudy (anonymous):

for (-2,0) and (2,8) m = (8-0)/(2-(-2)) = 8/4 = 2 now consider point(-2,0) we have y=mx+b hence we get -2 =2*0+b therefore b=-2 then equation in point form is y=2x-2

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

welcome